James decided to go and play basketball at the gym. James scored 40 and missed 10. What is the 95% confidence interval of the proportion of all shots he would have made.
Answer #1

n = 40 + 10

x = 40

p = 40/50 = 0.8

z = 1.96

The 95% confidence interval will be:

= p ± z*(√p(1-p)/n

= 0.8 ± 1.96*(√0.8(1-08)/50

= (0.6891, 0.9109)

The 95% confidence interval is between 0.6891 and 0.9109.

Please give me a thumbs-up if this helps you out. Thank you!